### v2

# The sequence of pseudo-equilibria describes the long-time behaviour of the NNLIF model with large delay

Abstract
There is a wide range of mathematical models that describe populations of large numbers of neurons. In this article, we focus on nonlinear noisy leaky integrate and fire (NNLIF) models that describe neuronal activity at the level of the membrane potential of neurons. We introduce a set of novel states, which we call “pseudo-equilibria”, and give evidence of their defining role in the behaviour of the NNLIF system when a significant synaptic delay is considered. The advantage is that these states are determined solely by the system’s parameters and are derived from a sequence of firing rates that result from solving a recurrence equation. We propose a new strategy to show convergence to an equilibrium for a weakly connected system with large transmission delay, based on following the sequence of pseudo-equilibria. Unlike with the direct entropy dissipation method, this technique allows us to see how a large delay favours convergence. We present a detailed numerical study to support our results. This study helps understand, among other phenomena, the appearance of periodic solutions in strongly inhibitory networks.
1 Introduction
Over the past few decades, a diverse range of research has emerged in the realm of partial differential equation (PDE) systems to model populations of large numbers of neurons. Depending on the variables taken into account to describe the activity of neurons, several families of PDE have been studied: Fokker-Planck equations including voltage and conductance variables [41, 8, 39, 40]; Fokker-Planck equations for uncoupled neurons [33, 34]; population density models of integrate and fire neurons with jumps [35, 26, 25, 24]; age structured equations (time elapsed models) [36, 37, 38] which are derived as mean-field limits of Hawkes processes [18, 17], McKean-Vlasov equations [1, 32], which are the mean-field equations of a large number of neurons described by the Fitzhugh-Nagumo equation [27], etc.
This article is devoted to the nonlinear noisy leaky integrate and fire (NNLIF) neuronal system, which is one of the simplest PDE models [30, 44, 35, 3, 2, 4]. We focus on its mesoscopic/macroscopic description through mean-field Fokker-Planck type equations [7, 16, 9, 12, 13, 11, 28, 42, 29], although it has also been studied at the microscopic level, using stochastic differential equations (SDE) [21, 20, 31]. Despite their simplicity, and the extensive scientific output devoted to them, fundamental questions remain about their long-term behaviour. The aim of this article is to contribute to the understanding of this issue. Specifically, we give evidence that the behaviour of NNLIF model with large delay is determined by a simple discrete system, which gives us a sequence of states, that we call “pseudo-equilibria”. The advantage of the discrete system lies in its simplicity: for instance, it allows for quick simulations that give accurate information about the NNLIF system. We will first describe the model more precisely, and then give our main results.
NNLIF systems describe the activity of large numbers of neurons, at the level of the membrane potential, defined as the potential difference between both sides of the neuron membrane. The PDE-based NNLIF family includes systems with different complexity depending on the neurophysiological properties that are taken into account [5]. In this article we shall focus on the following nonlinear Fokker-Planck equation [3, 20, 21, 31]:
$\partial_{t}p(v,t)+\partial_{v}\left[h\left(v,N(t-d)\right)p(v,t)\right]-a\partial^{2}_{v}p(v,t)=\delta(v-V_{R})N(t),$ (1.1)
which arises as the mean-field limit of a large set of $\mathcal{N}$ identical neurons which are connected to each other in a network, when $\mathcal{N}\to+\infty$. This PDE provides the evolution in time $t\geq 0$ of the probability density $p(v,t)>0$ of finding neurons at voltage $v\in\left(-\infty,V_{F}\right]$. A neuron spikes when its membrane voltage reaches the firing threshold value $V_{F}$, discharges immediately after, and has its membrane potential set back to the reset value $V_{R}$ ($V_{R}<V_{F}$), which is described by the right hand side of the equation and the boundary condition $p(V_{F},t)=0$. In addition, the model includes the delay $d$ in synaptic transmission. In this paper we assume the diffusion coefficient $a>0$ to be constant, and we always take the drift coefficient $h$ to be $h(v,N):=-v+bN$. This makes the system nonlinear since the firing rate $N=N(t)$ is computed as
$N(t):=-a\,\partial_{v}p(V_{F},t)\geq 0.$
The parameter $b$, called the connectivity parameter, shows how excitatory or inhibitory the network is. The case $b=0$ means that neurons are not connected with each other and the system becomes linear. Otherwise, if $b>0$ the network is average-excitatory, and if $b<0$ the network is average-inhibitory. Finally, assuming also $p(-\infty,t)=0$ the solution to the related Cauchy problem satisfies the conservation law:
[…]
for a given non-negative, integrable initial condition $p(v,0)=p_{0}(v)\geq 0$, with $N(t)=-a\partial_{v}p_{0}(V_{F})$ for $t\in[-d,0]$, and where $T>0$ is the maximal time of existence [16, 11, 42]. For the sake of simplicity, we consider that $\int_{-\infty}^{V_{F}}p_{0}(v)\,\mathrm{d}v=1$ and the diffusion coefficient $a=1$.
The probability stationary states, steady states, or equilibria, $p_{\infty}(v)$, of system (1.1) are non negative solutions to
$\left\{\begin{array}[]{l}\partial_{v}\left[\left(-v+bN_{\infty}\right)p_{\infty}(v)\right]-\partial^{2}_{v}p_{\infty}(v)=\delta(v-V_{R})N_{\infty},\\ N_{\infty}:=-\partial_{v}p_{\infty}(V_{F}),\ \text{and}\ \int_{-\infty}^{V_{F}}p_{\infty}(v)\,\mathrm{d}v=1,\end{array}\right.$ (1.2)
and are given by
$p_{\infty}(v)=N_{\infty}e^{-\frac{\left(v-bN_{\infty}\right)^{2}}{2}}\int_{\max(v,V_{R})}^{V_{F}}e^{\frac{\left(w-bN_{\infty}\right)^{2}}{2}}\,\mathrm{d}w.$ (1.3)
The stationary firing rate $N_{\infty}$ is implicitly given by the requirement of unit mass ($\int_{-\infty}^{V_{F}}p_{\infty}(v)\,\mathrm{d}v=1$), which is translated into the condition $N_{\infty}I(N_{\infty})=1$, where
$I(N):=\int_{-\infty}^{V_{F}}e^{-\frac{\left(v-bN\right)^{2}}{2}}\int_{\max(v,V_{R})}^{V_{F}}e^{\frac{\left(w-bN\right)^{2}}{2}}\,\mathrm{d}w\,\mathrm{d}v.$ (1.4)
Every stationary state corresponds to a solution of the implicit equation $N_{\infty}I(N_{\infty})=1$ for $N_{\infty}$. The number of steady states depends on the connectivity parameter $b$ [7]: for $b\leq 0$ (inhibitory case) there is only one, while for $b>0$ (excitatory case) there is more variety: there is only one if $b$ is small; there are no steady states if $b$ is large; and there are two for intermediate values of $b$ (the existence of at least two can be proved analytically; exactly two are observed numerically).
For the nonlinear Fokker-Planck Equation (1.1) there is a global in time solution if $d>0$ [11, 42], and, if $d=0$ and $b<0$ (the average-inhibitory case) [16]. However, for the average-excitatory case ($b>0$), blow-up of the solution may occur, and in this case the maximal time of existence is given by the first time at which the firing rate $N(t)$ diverges [16, 7] if there is no transmission delay and the initial condition is sufficiently concentrated around the threshold potential $V_{F}$. The blow-up is avoided if some transmission delay or stochastic discharge potential are considered [9]. The analogous criteria for existence and blow-up phenomena were studied in [21], for the associated microscopic system. Moreover, the notion of physical solutions to SDE was given in [20] and the authors proved physical solutions are global on time, solutions continue after system synchronization, this is after the blow-up phenomenon. Understanding what “physical solution” may mean for the Fokker-Planck equation is an open problem. It was numerically analysed in [10], and behaviour after the explosion was studied in [22] by a time change of variable which dilates the blow-up time.
Let us give a brief review of the existing results on asymptotic behaviour. Most of the literature on asymptotic behaviour of the model (1.1) is based on the entropy dissipation method (see [7, 9, 15, 12, 11]), considering a standard relative entropy functional to estimate the distance between the solution $p$ and the stationary solution $p_{\infty}$. This strategy has given results only for weakly connected networks (small values of the connectivity parameter $b$), that is, almost linear systems. Recently, approaches based on Doeblin & Harris’s theorems in probability have been used in the study of the asymptotic behaviour of various equations related to neural networks, for both elapsed time [14] and integrate-and-fire models [24, 43]. Again, both strategies (entropy method and Harris-type theorems) seem to be suitable only for nearly linear systems. Results on the long-term behaviour for general connectivity have been recently given in [6] by using a new strategy based on the spectral gap properties of the linearised equation. On the other hand, [11, 42] proved that there are no periodic solutions if $b$ is large enough, $V_{F}\leq 0$ and a transmission delay is considered. Recently in [29] it was shown that periodic solutions can appear in the average-inhibitory case if a large delay is considered for an approximation of the integrate-and-fire model. Moreover, for the more realistic model (considering as different populations the excitatory and inhibitory neurons, or neurons with refractory periods) and the stochastic discharge potential model, the results in [13, 9, 28, 23, 45] numerically show periodic solutions.
Let us describe the main ideas in the present paper. We wish to examine a discrete sequence determined by the parameters of the nonlinear system (1.1). For a given $N\in\mathbb{R}^{+}$ we define the pseudo-equilibrium associated to $N$ as
$p_{\tiny\mbox{pseudo}}(v):=\tilde{N}e^{-\frac{\left(v-bN\right)^{2}}{2}}\int_{\max(v,V_{R})}^{V_{F}}e^{\frac{\left(w-bN\right)^{2}}{2}}\,\mathrm{d}w,\quad\mbox{with}\quad\tilde{N}=\frac{1}{I(N)}.$ (1.5)
This is just an equilibrium for the linear version of equation (1.1), where the dependence of $h$ on $N(t-d)$ is “frozen” at a certain value $N$. The constant $\tilde{N}$ guarantees unit mass: $\int_{-\infty}^{V_{F}}p_{\tiny\mbox{pseudo}}(v)\,\mathrm{d}v=1$. The profile $p_{\tiny\mbox{pseudo}}$ is an equilibrium of the NNLIF system (1.1) if and only if and only if $\tilde{N}=N$, that is, if and only if $NI(N)=1$.
Our guiding idea is that for a large delay $d$, the NNLIF system (1.1) behaves almost as a linear system in time intervals of length $d$. Assume that we start with a constant initial condition on $[-d,0]$, which has a firing rate $N_{0,\infty}$. Then the NNLIF system is exactly linear on $[0,d]$; is $d$ is large, we expect the solution at $t=d$ to be close to the linear equilibrium associated to a fixed firing rate $N_{0,\infty}$, that is, the pseudo-equilibrium associated to a certain firing rate $N_{1,\infty}$. This corresponds to taking $N=N_{0,\infty}$ and $\tilde{N}=N_{1,\infty}$ in (1.5). On the time interval $[d,2d]$, if on $[0,d]$ the firing rate has remained close to $N_{1,\infty}$, we expect the solution to again behave linearly, and approach another pseudo-equilibrium with firing rate $N_{2,\infty}$. Iterating this we define a sequence of pseudo-equilibria determined solely by the initial value $N_{0,\infty}$ (and the given parameters $V_{R}$, $V_{F}$, $b$). We are not able to rigorously prove that this sequence represents the behaviour of the system with long delays except for the case of small $b$ (see Section 3). However, we present strong numerical evidence that this sequence does represent the overall behaviour of the system (cf. Section 4). In particular, this sequence shows periodic behaviour for $b$ negative enough; converges to the stable equilibrium of (1.1) in the region where solutions (1.1) do the same; and approaches a plateau state when solutions to (1.1) are expected to do this.
Let us describe this plateau state more closely, since it is also an important evidence for the usefulness of the pseudo-equilibria sequence. Recently, in [10] we numerically observed the formation of a new profile, which describes networks with uniformly distributed membrane potentials between the reset value $V_{R}$ and the threshold value $V_{F}$. We called these states “plateau” distributions. These profiles appear in two situations: 1) strongly connected systems (high connectivity parameter $b$) with synaptic delay $d>0$, and 2) systems with $b=V_{F}-V_{R}$, without delay ($d=0$) or with very small transmission delay value. Numerically we observe that plateau states appear to be related to pseudo-equilibria in the following way: Figure 1 shows the comparison of the plateau distributions found for $b=1.5$ (left plot) and $b=2.2$ (right plot), with delay transmission, and several pseudo-equilibria. There is a high coincidence between the plateau distributions and the profiles (1.5) as $N$ increases. Furthermore, on the way to the plateau state, the dynamics of the particle system appears to pass through pseudo-equilibria of the form (1.5) as $N$ increases. For $b=1.5$ the system (1.1) has two steady states, and it can evolve towards a plateau distribution (shown on the left plot) if the initial condition is far from the stationary state with lowest firing rate. For $b=2.2$ there are no stationary solutions and the system evolves towards a plateau distribution.
Figure 1: Comparison of “plateau” distributions with profiles (1.5). Graphs taken from [10]. Left: $b=1.5$. Right: $b=2.2$.
This behaviour suggests that we may think of the pseudo-equilibria sequence as a tool to analyse the asymptotic behaviour of the system when a large synaptic delay is taken into account. The aims of this work are twofold:
  1. 1.
    to analyse the sequence of pseudo-equilibria determined by the associated discrete system, and
  2. 2.
    to establish a connection between the long-term behaviour of the pseudo-equilibria sequence and the solutions of the nonlinear system (1.1).
In Section 2 we define and analyse the firing rate and the pseudo-equilibria sequences. In Section 3 we prove the exponential convergence to equilibrium for the nonlinear system (1.1) for small connectivity parameter $b$ and large delay $d$, by “following” the sequence of pseudo-equilibria, under some technical assumptions on the linear systems. A similar result was previously established through the use of the entropy method, and we present present here different strategy which considers the relative entropy to the sequence of pseudo-equilibria. This approach looks promising to study a broader range of phenomena (such as periodic solutions or the approach of the plateau state), and served as a catalyst for the main ideas in our paper [6], where we prove local convergence results depending on the values of the transmission delay and the connectivity parameter. In Section 4 we provide the numerical evidence to support the connection between the sequence of pseudo-equilibria studied in Section 2 with the nonlinear system. This numerical work also suggests several conjectures on the possible extension of the results of Section 3 to cases with an arbitrary value of the parameter $b$. Section 5 is devoted to conclusions, discussion and possible extensions of our work. We also include an appendix with some of the technical results needed to prove Theorem 2.5.
2 Firing rate and pseudo-equilibria sequences
This section deals with the study of sequences of firing rates, $\left\{N_{k,\infty}\right\}_{k\geq 0}$, and pseudo-equilibria, $\left\{p_{k,\infty}\right\}_{k\geq 0}$. Let us first give their definition:
Definition 2.1 (Firing rate sequence).
Given $b\in\mathbb{R}$, $V_{R}<V_{F}\in\mathbb{R}$, and $I$ the function (1.4), the firing rate sequence $\left\{N_{k,\infty}\right\}_{k\geq 0}$ associated to an initial firing rate $N_{0,\infty}>0$ is recursively defined by
$N_{k+1,\infty}:=\frac{1}{I(N_{k,\infty})}\qquad k=0,1,2,\dots$ (2.1)
This sequence is well defined since $I(N)>0$ for all $N>0$. Associated to a firing rate sequence $\{N_{k,\infty}\}_{k\geq 0}$ we define the following pseudo-equilibria sequence:
Definition 2.2 (Pseudo-equilibria sequence).
Given a firing rate sequence $\left\{N_{k,\infty}\right\}_{k\geq 0}$ its associated pseudo-equilibria sequence is the sequence
[…]
The behaviour of these sequences depends only on the values of the connectivity parameter $b$, the diffusion coefficient $a$ (which we consider to be 1), the reset potential $V_{R}$ and the threshold potential $V_{F}$. Their definition is completely detached from the dynamics of the nonlinear system (1.1).
In the following theorems we show the monotonicity of the firing rate sequence $\left\{N_{k,\infty}\right\}_{k\geq 0}$, in terms of the number of solutions to the implicit equation
$NI(N)=1,\ \mbox{with}\ 0\leq N.$ (2.3)
We have denoted the unknown in Equation (2.3) by $N$ and hope that it will not cause any confusion with the firing rate of the nonlinear system (1.1). In this section we never consider the system (1.1), since we are only concerned with the behaviour of the firing rate sequence. The number of solutions to Equation (2.3) was studied in [7]. Analytically it was shown that for $b<0$ or for a small positive value of $b$ there is only one solution, there is no solution if $b>0$ is large, and there are at least two for intermediate positive values of $b$. The proof is based on the study of the monotone function $1/I(N)$ (increasing if $b>0$ and decreasing if $b<0$). However, rigorously proving the exact number of solutions in the excitatory case ($b>0$) is not easy, due to the complexity of $I$. Numerically we observe that for positive $b$ there are a maximum of two solutions, and at the point where the equation switches from having two solutions to none the function $1/I(N)$ has slope equal to 1 and positive second derivative (see right plot of Figure 2). These are the scenarios we consider in the following theorem for the case $b>0$:
Theorem 2.3 (Monotonicity of the firing rate sequence $\left\{N_{k,\infty}\right\}_{k\geq 0}$ for excitatory networks).
Let us consider $0<b$ and a firing rate sequence $\left\{N_{k,\infty}\right\}_{k\geq 0}$ (as given by Definition 2.1).
  1. 1.
    If $N^{*}$ is the only solution to Equation (2.3), then:
    1. (a)
      If $N\to 1/I(N)$ crosses the line $N\to N$ at $N=N^{*}$ ($N<1/I(N)$ if $N<N^{*}$ and $1/I(N)<N$ if $N^{*}<N$):
      • If $N^{*}\leq N_{0,\infty}$ then $\left\{N_{k,\infty}\right\}_{k\geq 0}$ is a decreasing sequence tending to $N^{*}$.
      • If $N_{0,\infty}\leq N^{*}$ then $\left\{N_{k,\infty}\right\}_{k\geq 0}$ is an increasing sequence tending to $N^{*}$.
    2. (b)
      If $N\to 1/I(N)$ cut tangentially the line $N\to N$ at $N=N^{*}$ being $N<1/I(N)$ if $N\neq N^{*}$:
      • If $N^{*}\leq N_{0,\infty}$ then $\left\{N_{k,\infty}\right\}_{k\geq 0}$ diverges.
      • If $N_{0,\infty}\leq N^{*}$ then $\left\{N_{k,\infty}\right\}_{k\geq 0}$ is an increasing sequence tending to $N^{*}$.
  2. 2.
    If Equation (2.3) has no solutions, then $\left\{N_{k,\infty}\right\}_{k\geq 0}$ diverges.
  3. 3.
    If Equation (2.3) has two solutions: $N_{1}^{*}$ and $N_{2}^{*}$ ($N_{1}^{*}<N_{2}^{*}$), then:
    • If $N_{0,\infty}\leq N_{1}^{*}$ then $\left\{N_{k,\infty}\right\}_{k\geq 0}$ is an increasing sequence which tends to $N_{1}^{*}$.
    • If $N_{1}^{*}\leq N_{0,\infty}\leq N_{2}^{*}$ then $\left\{N_{k,\infty}\right\}_{k\geq 0}$ is a decreasing sequence which tends to $N_{1}^{*}$.
    • If $N_{2}^{*}\leq N_{0,\infty}$ then $\left\{N_{k,\infty}\right\}_{k\geq 0}$ diverges.
Proof.
The function $I$ (see (1.4)) is a $C^{\infty}(0,\infty)$ function, which was studied in the proof of [7, Theorem 3.1], and can be rewritten as $I(N)=\int_{0}^{\infty}e^{-s^{2}/2}e^{-sbN}\frac{e^{sV_{F}}-e^{sV_{R}}}{s}\ \,\mathrm{d}s.$ Its $k$-th order derivative is
$I^{(k)}(N)=(-b)^{k}\int_{0}^{\infty}s^{k-1}e^{-s^{2}/2}e^{-sbN}\left(e^{sV_{F}}-e^{sV_{R}}\right)\ \,\mathrm{d}s,$ (2.4)
and for $b$ positive:
  • $\frac{1}{I(N)}$ is an increasing function.
  • $\frac{1}{I(0)}<\infty$.
  • $\lim_{N\to\infty}\frac{1}{I(N)}=\infty$.
The firing rate sequence is a solution to the recursive equation
$N_{k+1,\infty}=f(N_{k,\infty}),\quad f:\mathbb{R}_{0}^{+}\to\mathbb{R}^{+},\ f(x):=\frac{1}{I(x)},$ (2.5)
and we have seen that $f$ is an increasing function. The behaviour of this type of discrete system is well known: its solutions are monotonic, and either diverge or converge to an equilibrium of the system. With the hypotheses of the theorem we have the following possibilities, since $f(0)>0$:
  • If $N^{*}$ is the only solution to Equation (2.3) then:
    • If $f(N)$ crosses the line $N\to N$ at $N=N^{*}$.
        $N<f(N)$ if $N<N^{*}$.
        $f(N)<N$ if $N^{*}<N$.
    • If $f(N)$ cut tangentially the line $N\to N$ at $N=N^{*}$ being $N<f(N)$ if $N\neq N^{*}$.
        $N<f(N)$ if $N\neq N^{*}$.
  • If Equation (2.3) has two solutions, $N_{1}^{*}$ and $N_{2}^{*}$, with $N_{1}^{*}<N_{2}^{*}$, then:
    • $N<f(N)$ if $N\in[0,N_{1}^{*})\cup(N_{2}^{*},+\infty)$.
    • $f(N)<N$ if $N\in(N_{1}^{*},N_{2}^{*})$.
  • If Equation (2.3) has no solutions, then $N<f(N)$ for all $N\geq 0$, because $0<f(0)$.
The proof is easily done by analysing the monotonicity of the sequence in each interval bounded by equilibria.
  1. 1.
    If the initial condition is taken in an interval in which $N<f(N)$ then the sequence is increasing, because $N_{k,\infty}<f(N_{k,\infty})=N_{k+1,\infty}$.
  2. 2.
    If the initial condition is taken in an interval in which $f(N)<N$ then the sequence is decreasing, because $N_{k+1,\infty}=f(N_{k,\infty})<N_{k,\infty}$.
Whenever the sequence is decreasing, it is bounded below by a constant solution of Equation (2.5) and therefore converges to it (because there are no more equilibria in that interval). If the sequence is increasing, it either converges to an equilibrium, if bounded, or diverges. ∎
Remark 2.4.
As a consequence of the convergence of the firing rate sequences, given in Theorem 2.3, it follows:
  1. 1.
    If $N^{*}$ is the only solution to Equation (2.3), then:
    1. (a)
      In case 1.(a) of Theorem 2.3, $\frac{d}{dN}\frac{1}{I(N)}|_{N^{*}}<1$.
    2. (b)
      In case 1.(b) of Theorem 2.3, $\frac{d}{dN}\frac{1}{I(N)}|_{N^{*}}=1$.
    This behaviour is shown in right plot of Figure 2 with $b=0$ or $b=0.5$ for the first case and $b=2.1$ the second case.
  2. 2.
    If Equation (2.3) has two solutions: $N_{1}^{*}$ and $N_{2}^{*}$ ($N_{1}^{*}<N_{2}^{*}$), then:
    • $0\leq\frac{d}{dN}\frac{1}{I(N)}|_{N_{1}^{*}}\leq 1$.
    • $1<\frac{d}{dN}\frac{1}{I(N)}|_{N_{2}^{*}}$.
    This behaviour is shown in right plot of Figure 2 with $b=1.5$.
We analyse the inhibitory case ($b<0$) in the following theorem, where two different behaviours appear although in this case Equation (2.3) has only one solution.
Theorem 2.5 (Monotonicity of the firing rate sequence $\left\{N_{k,\infty}\right\}_{k\geq 0}$ for inhibitory networks).
Assume $b<0$ and consider a firing rate sequence $\left\{N_{k,\infty}\right\}_{k\geq 0}$. Then there exists a value $b^{*}<b$ of the connectivity parameter $b$ such that:
  • If $b^{*}<b<0$ the sequence $\left\{N_{k,\infty}\right\}_{k\geq 0}$ tends to the unique solution to Equation (2.3), $N^{*}$.
  • If $b<b^{*}$, there exist two values $N^{-}$, $N^{+}$, $0\leq\min(N^{-},N^{+})<N^{*}<\max(N^{-},N^{+})$, such that the sequence $\left\{N_{k,\infty}\right\}_{k\geq 0}$ tends to the 2-cycle $\{N^{-},N^{+}\}$.
Proof.
The proof, as in Theorem 2.3, is based on the study of the recursive equation:
$N_{k+1,\infty}=f(N_{k,\infty}),\quad\ f(x):=\frac{1}{I(x)},$ (2.6)
where the function $I$, defined in (1.4), is a $C^{\infty}(0,\infty)$ increasing function if $b<0$ (see (2.4) with $k=1$). Therefore, for the inhibitory case, $f:\left[0,\frac{1}{I(0)}\right]\to\left[0,\frac{1}{I(0)}\right]$ since the function $1/I$ is decreasing, and, with the possible exception of the initial datum, $N_{0,\infty}$, all the other terms in the sequence fall within the interval $\left[0,\frac{1}{I(0)}\right]$. Moreover, in this case, Equation (2.3) has an unique solution, $N^{*}$, because $1/I(N)$ is a decreasing continuous function, which tends towards 0 and $0<1/I(0)$. The behaviour of this type of discrete system is easy to study, because $f$ is an decreasing function and the solutions are bounded; they tend to the unique equilibrium or towards 2-cycle. The proof of this result is as follows: we denote by $N^{*}$ the equilibrium of (2.6) and suppose $N_{0,\infty}\neq N_{2,\infty}$ (otherwise, the solution would be a 2 cycle, which would be trivial if $N_{0,\infty}=N_{2,\infty}=N^{*}$). Thus, using that $f$ is a decreasing function and $F:=f\circ f$ is increasing:
  • If $N_{0,\infty}\leq N_{2,\infty}$, then $N_{3,\infty}=f(N_{2,\infty})\leq N_{1,\infty}=f(N_{0,\infty})$ and $N_{2,\infty}=F(N_{0,\infty})\leq N_{4,\infty}=F(N_{2,\infty})$. Therefore, by induction we prove that $\{N_{2k,\infty}\}_{\{k\geq 0\}}$ is an increasing sequence and $\{N_{2k+1,\infty}\}_{\{k\geq 0\}}$ is a decreasing sequence.
  • If $N_{0,\infty}\geq N_{2,\infty}$, then $N_{3,\infty}=f(N_{2,\infty})\geq N_{1,\infty}=f(N_{0,\infty})$ and $N_{2,\infty}=F(N_{0,\infty})\geq N_{4,\infty}=F(N_{1,\infty})$. Therefore, by induction we prove that $\{N_{2k,\infty}\}_{\{k\geq 0\}}$ is a decreasing sequence and $\{N_{2k+1,\infty}\}_{\{k\geq 0\}}$ is an increasing sequence.
Moreover:
  • If $N_{0,\infty}\leq N^{*}$, then $N^{*}=f(N^{*})\leq f(N_{0,\infty})=N_{1,\infty}$ and $N_{2,\infty}=f(N_{1,\infty})\leq f(N^{*})=N^{*}$. Therefore, by induction we prove that $0\leq N_{2k,\infty}\leq N^{*}\leq N_{2k+1,\infty}\leq\frac{1}{I(0)}$ for $k=0,1,2,\ldots$.
  • If $N^{*}\leq N_{0,\infty}$, then $f(N_{0,\infty})=N_{1,\infty}\leq N^{*}=f(N^{*})$ and $f(N^{*})=N^{*}\leq N_{2,\infty}=f(N_{1,\infty})$. Therefore, by induction we prove that $0\leq N_{2k+1,\infty}\leq N^{*}\leq N_{2k,\infty}\leq\frac{1}{I(0)}$ for $k=0,1,2,\ldots$.
So that in either case, both sequences, $\{N_{2k,\infty}\}_{\{k\geq 0\}}$ and $\{N_{2k+1,\infty}\}_{\{k\geq 0\}}$, are monotonic and bounded, hence convergent. Thus, considering $N^{-}:=\lim_{k\to\infty}N_{2k+1,\infty}$ and $N^{+}:=\lim_{k\to\infty}N_{2k,\infty}$ we obtain, since $f$ is a continuous function, that $f(N^{+})=N^{-}$ and there are two possibilities:
  1. 1.
    $N^{-}=N^{+}=N^{*}$, so that the system tends towards the equilibrium.
  2. 2.
    $N^{-}\neq N^{+}$, so that the system tends towards the 2-cycle $\{N^{-},N^{+}\}$, where $\min(N^{-},N^{+})<N^{*}<\max(N^{-},N^{+})$.
[…]
$\frac{dN_{b}^{*}}{db}=\frac{-N\partial_{b}I(b,N_{b}^{*})}{I(b,N_{b}^{*})+N_{b}^{*}\partial_{N}I(b,N_{b}^{*})}\geq 0.$
Consequently, $N^{*}_{b}$ is an increasing function bounded from below by 0, so it has a limit when $b$ tends to $\infty$- ∞. On the other hand, when $b$ tends to $\infty$- ∞, $\frac{1}{I(b,N)}$ loses continuity at $N=0$, because $\frac{1}{I(b,0)}>0$, while it vanishes for $0<N<1/I(b,0)$. Therefore, $\overline{N}^{*}:=\lim_{b\to-\infty}N^{*}_{b}=0$, because, if that limit is positive, we get a contradiction, since $00=\lim_{b\to-\infty}\frac{1}{I(b,\overline{N}^{*})}=\lim_{b\to-\infty}\frac{1}{I(b,N_{b}^{*})}=\lim_{b\to-\infty}N_{b}^{*}=\overline{N}^{*}>$ 0 (this limit was also proven in [29][Lemma 2.2]). Thus, in a certain sense, we can understand that the system loses its equilibrium, when $b$ tends to $\infty$- ∞, due to loss of continuity of function $1/I$.
To determine whether or not the equilibrium of Equation (2.6), $N^{*}_{b}$, is asymptotically stable, we check if $|f^{\prime}(N^{*}_{b})|$ is less than or greater than 1. With this aim, we define
$g(b):=f^{\prime}(N^{*}_{b})=\frac{d}{dN}\frac{1}{I(N)}|_{N^{*}_{b}}=\frac{-\partial_{N}I(b,N^{*}_{b})}{I(b,N^{*}_{b})^{2}}=-{N_{b}^{*}}^{2}\partial_{N}I(b,N^{*}_{b}),$
which is a continuous increasing function in $(-\infty,0]$, due to the positivity of $g^{\prime}(b)$, as proved in Lemma A.1. We note that $\lim_{b\to-\infty}g(b)=-\infty$ and $g(0)=f^{\prime}(N^{*}_{0})=0$, thus there exists $b^{*}$ such that $g(b^{*})=-1$. Then, if $b<b^{*}$, thus $g(b)<-1$, which means that $N_{b}^{*}$ is unstable, while for $b>b^{*}$, $g(b)>-1$ and $N_{b}^{*}$ is stable. And from what has been proven above, if the equilibrium is unstable, the system tends to a 2-cycle. ∎
Remark 2.6.
Two details could not be proved in Theorem 2.5, due to the complexity of the function $I$: the stability of the equilibrium in the critical case, $b=b^{*}$, and the uniqueness of the two cycles, in the case $b<b^{*}$. However, we can check both numerically. Figure 3 shows the function $F(N)=f\circ f(N)$, for cases with $b<b^{*}$. $F$ has only $3$ fixed points, i.e., the equilibrium $N^{*}$ and cycle $\left\{{N^{-},N^{+}}\right\}$ corresponding to each value of $b$. Moreover, the 2-cycles appear as a bifurcation of the equilibrium, which is asymptotically stable for $b=b^{*}$, as this value is the starting point for the formation of 2-cycles. We also note that the 2-cycle $\left\{{N^{-},N^{+}}\right\}$ becomes $\left\{0,\frac{1}{I(0)}\right\}$, approximately $\left\{0,0.12\right\}$, when $b$ tends to $\infty$- ∞, as $N_{b}^{*}$ tends to 0. For $b>b^{*}$ we observe an unique fixed point for $F(N)$, corresponding with the equilibrium $N^{*}$.
Remark 2.7.
As a consequence of the convergence of the firing rate sequences, given in Theorem 2.5, it follows:
  1. 1.
    If $b_{*}<b<0$, then $-1\leq\frac{d}{dN}\frac{1}{I(N)}|_{N^{*}}<0$, for $N^{*}$, the only solution to Equation (2.3).
  2. 2.
    If $b<b_{*}$, then $\frac{d}{dN}\frac{1}{I(N)}|_{N^{*}}\leq-1$, for $N^{*}$, the only solution to Equation (2.3).
We have estimated numerically $b^{*}$ around $-9.4$ (see Figure 3), the left plot of Figure 2 shows the function $\frac{1}{I(N)}$ for $b<-9.4$ and for $b>-9.4$, where we can observe its slope at $N^{*}$.
Figure 2: Function $\boldsymbol{\frac{1}{I(N)}}$ (see (1.4)) for different values of the connectivity parameter $\boldsymbol{b}$.
Figure 3: Function $\boldsymbol{F:=f\circ f}$ with $\boldsymbol{f=1/I(N)}$ (see (1.4)) for different values of the connectivity parameter $\boldsymbol{b}$.
We recall that the number of solutions to the Equation (2.3) gives us the number of steady states of the nonlinear system (1.1). And that number depends on the value of $b$ (see [7]). For the excitatory case ($b>0$) there is only one steady state if $b\leq V_{F}-V_{R}$, if $V_{F}-V_{R}<b$ there are two possibilities: there are at least two steady states (numerically no more than two steady states have been observed), or, if $b$ is large enough, there is not steady states. The case $b=V_{F}-V_{R}$ is very interesting because this is the limit value for which the notion of physical solutions makes sense. Numerically this case was analysed in [10] and without delay ($d=0$) the particle system evolves to a plateau distribution, if the initial condition is concentrated enough around $V_{F}$. In a limit sense, this plateau profile is a stationary state of the system for $b=V_{F}-V_{R}$. However, the system tends to the unique steady state with bounded firing rate, if a large transmission delay is considered. For the inhibitory case ($b<0$) there is only one steady state. Therefore, Theorems 2.3 and 2.5 can also be read in terms of the number of the steady states of the nonlinear system (1.1).
Remark 2.8.
[…]